OpenStudy (j.williams4):

1. Which ray is a bisector of

OpenStudy (anonymous):

1. BD

OpenStudy (mathstudent55):

You should ask only one question per post. Also, I think you need figures for these questions.

OpenStudy (mathstudent55):

1. The question is about an angle bisector. Do you know what an angle bisector is?

OpenStudy (anonymous):

2. 3x = 5x - 10 2x = 10 x = 5 So : 5x - 10 = 5(5) - 10 = 15.

OpenStudy (anonymous):

3. 10y = 8y + 4 2y = 4 y = 2

OpenStudy (anonymous):

4. Angle(DBE) = Twice of (8y+4) = 2*(8*2 + 4) = 2*20 = 40

OpenStudy (anonymous):

5. 7x + 6y = 7*5 + 6*2 = 35 + 12 = 47 (: